OpenStudy (anonymous):

Hi, is there any specific formula or series for tan(nx) in terms of n and tan(x)?

OpenStudy (anonymous):

salam

OpenStudy (anonymous):

salam!

OpenStudy (anonymous):

ye negahi be in bendaz http://mathhelpforum.com/trigonometry/177739-tan-x.html

OpenStudy (anonymous):

aaaali bud! mamnoon!

OpenStudy (anonymous):

xahesh mikonam...baram jalebe age beshe formoole kolli bara n be dast ovord...

OpenStudy (anonymous):

hatman roosh kar mikonam

mathslover (mathslover):

nikasi thadi nake? putisi khudi khudi chinka phate pe ..

OpenStudy (anonymous):

lol

mathslover (mathslover):

kitcha "lol" phika jhinka jhinka tooon..

Parth (parthkohli):

@ParsaN Salam, zahesh mikonam Ka Sep... @mukushla "mathematical formula". zey.

mathslover (mathslover):

translate.google.com worked @ParthKohli :P

OpenStudy (anonymous):

:D

Parth (parthkohli):

@mathslover What's that? That's not the Irani I am learning at school...

OpenStudy (anonymous):

that makes sense parth :D

mathslover (mathslover):

:D

OpenStudy (anonymous):

\[\cos nx+i \sin nx=e^{inx}=(e^{ix})^n=(\cos x+i \sin x)^n\\=\sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right)i^k \ \sin^k x \ \cos^{n-k}x=\cos^n x\sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right)i^k \ \sin^k x \ \cos^{-k}x\\=\cos^n x\sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right)i^k \ \tan^k x \]so we can say \[\tan nx=\frac{Im\left( \sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right)i^k \ \tan^k x \right)}{Re\left( \sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right)i^k \ \tan^k x \right)}\]

OpenStudy (anonymous):

and finally\[\tan nx=\frac{n\tan x-\left(\begin{matrix}n \\ 3\end{matrix}\right)\tan^3 x+\left(\begin{matrix}n \\ 5\end{matrix}\right)\tan^5 x+...}{1-\left(\begin{matrix}n \\ 2\end{matrix}\right)\tan^2 x+\left(\begin{matrix}n \\ 4\end{matrix}\right)\tan^4 x+...}\]

OpenStudy (anonymous):

agha ino man takmil kardam..khedmate shoma..\[\large \tan nx=\frac{ \sum_{k=0}^{\lfloor \frac{n-1}{2} \rfloor} \left(\begin{matrix}n \\ 2k+1\end{matrix}\right) \ \tan^{2k+1} x}{\sum_{k=0}^{\lfloor \frac{n}{2} \rfloor} \left(\begin{matrix}n \\ 2k\end{matrix}\right) \ \tan^{2k} x }\]

OpenStudy (anonymous):

fantastic! thanks a lot @mukushla ...

OpenStudy (anonymous):

:)